# pEBR: A Probabilistic Approach to Embedding Based Retrieval

### v2

@@ 6 Limitations
While our probabilistic model demonstrates significant improvements in retrieval accuracy and system performance, it remains fundamentally constrained by the two-tower and late-interaction retrieval paradigms (e.g., DSSM architectures). Notably, the emerging paradigm of generative retrieval Kuai et al. (2024); Li et al. (2024); Deng et al. (2025) using large language models (LLMs) presents both opportunities and challenges that our current framework does not address. Three key directions for future research emerge: 1) Developing hybrid architectures that combine the reliability of our probabilistic scoring with the semantic flexibility of generative retrieval, particularly for dynamic cutoff optimization. 2) Investigating constrained generation techniques to maintain the precision advantages of our approach while reducing LLM hallucination rates by adaptive distribution.
@@ 7 Acknowledgments
We would like to express our sincere gratitude to Lin Liu, Songlin Wang, and Sulong Xu for their valuable guidance and technical support throughout this research. Their insightful suggestions and continuous encouragement were instrumental in the successful completion of this work.

### v3

@@ 6 Limitations
While our probabilistic model demonstrates significant improvements in retrieval accuracy and system performance, it remains inherently constrained by the two-tower and late-interaction retrieval paradigms (e.g., DSSM architectures). Notably, the emerging paradigm of generative retrieval Kuai et al. (2024); Li et al. (2024); Deng et al. (2025) powered by large language models (LLMs) introduce new opportunities and challenges that our current framework does not address. We identify two promising directions for future research: 1) developing hybrid architectures that integrate the robustness and reliability of our probabilistic modeling with the semantic flexibility of generative retrieval, particularly for dynamic cutoff optimization. 2) Exploring constrained generation techniques that retain the precision and advantages of our current approach while mitigating LLM hallucinations through adaptive distribution modeling, thereby enhancing both accuracy and reliability in retrieval tasks.
@@ 7 Acknowledgments
We would like to express our sincere gratitude to Lin Liu, Songlin Wang, and Sulong Xu for their valuable guidance and technical support throughout this research. Their insightful suggestions and continuous encouragement were instrumental in shaping the direction of our work.